DK4L:

Compute (6 + 4i)^2(6+4i) ​2 ​​ -20 + 48i 20 + 48i 40 + 48i -40 + 48i

100kpimpcess:

@Hoodmemes and @supie

DK4L:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @100kpimpcess @Hoodmemes and @supie \(\color{#0cbb34}{\text{End of Quote}}\) ok thnx

100kpimpcess:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @DK4L \(\color{#0cbb34}{\text{Originally Posted by}}\) @100kpimpcess @Hoodmemes and @supie \(\color{#0cbb34}{\text{End of Quote}}\) ok thnx \(\color{#0cbb34}{\text{End of Quote}}\) mhm no problem

DK4L:

Does anybody know how to do this?

100kpimpcess:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @DK4L Does anybody know how to do this? \(\color{#0cbb34}{\text{End of Quote}}\) no

100kpimpcess:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @100kpimpcess \(\color{#0cbb34}{\text{Originally Posted by}}\) @DK4L Does anybody know how to do this? \(\color{#0cbb34}{\text{End of Quote}}\) no \(\color{#0cbb34}{\text{End of Quote}}\) holon wait

DK4L:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @100kpimpcess \(\color{#0cbb34}{\text{Originally Posted by}}\) @100kpimpcess \(\color{#0cbb34}{\text{Originally Posted by}}\) @DK4L Does anybody know how to do this? \(\color{#0cbb34}{\text{End of Quote}}\) no \(\color{#0cbb34}{\text{End of Quote}}\) holon wait \(\color{#0cbb34}{\text{End of Quote}}\) ight

supie:

Why does it have a `2` under the problem?

DK4L:

(6+4i)^2

supie:

Ohhh ok

100kpimpcess:

i dont know tbh wut grade u in

DK4L:

10th

100kpimpcess:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @DK4L 10th \(\color{#0cbb34}{\text{End of Quote}}\) yuh im able to do some 10th grad work but not this im in 8th

DK4L:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @100kpimpcess \(\color{#0cbb34}{\text{Originally Posted by}}\) @DK4L 10th \(\color{#0cbb34}{\text{End of Quote}}\) yuh im able to do some 10th grad work but not this im in 8th \(\color{#0cbb34}{\text{End of Quote}}\) ight thats ok

darkknight:

Compute (6 + 4i)^2(6+4i)^2 ​ ​​ -20 + 48i 20 + 48i 40 + 48i -40 + 48i 36+16i^2+24i)^2 = -176 + 960i Wghich is not an answer choice so what is ur q, reformat for clarity plz

supie:

"Wghich"?

supie:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @darkknight Compute (6 + 4i)^2(6+4i)^2 ​ ​​ -20 + 48i 20 + 48i 40 + 48i -40 + 48i 36+16i^2+24i)^2 = -176 + 960i Wghich is not an answer choice so what is ur q, reformat for clarity plz \(\color{#0cbb34}{\text{End of Quote}}\) The post is closed so he doesn't need to "reformat for clarity"

darkknight:

@supie chill dawg he still might not have his answer

supie:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @darkknight @supie chill dawg he still might not have his answer \(\color{#0cbb34}{\text{End of Quote}}\) If he didn't have the answer he would just re-post it,

DK4L:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @darkknight @supie chill dawg he still might not have his answer \(\color{#0cbb34}{\text{End of Quote}}\) I have the anwser but thx for ur help/and its not he its she

100kpimpcess:

hmm